# Understanding and Mitigating the Impacts of Differentially Private Census Data on State Level Redistricting

### v1

@@ Acknowledgements
We thank Micah Altman, John Abowd, and Moon Duchin for valuable feedback. CC and AC were supported by the DARPA SIEVE program under Agreement No. HR00112020021. Preliminary work on this project was supported by the Boston University Center for Antiracist Research.
Appendix A Convergence and Robustness Tests
@@
To provide evidence of the convergence of our chains, we report the Gelman-Rubin split-$\hat{R}$ metric and the effective sample size (ESS) for each legislative geography we study in Table 3. Traditionally, an $\hat{R}$ of below 1.1 has been accepted as evidence of convergence, although more recent work recommends that using a threshold of 1.01 for $\hat{R}$ along with a rank-normalized ESS of at least 400 for more robust results. We test for the convergence of our measurement of population balance (i.e. whether our estimate of the mean of the function $f(\mathcal{P})=\mathds{1}\left[\mathsf{dev_{swap}}(\mathcal{P})>0.1\right]$ has converged, where $\mathcal{P}$ is a plan sampled from the stationary distribution of our Markov chain) and our measurement of MMD discrepancy (i.e. $f(\mathcal{P})=\mathsf{MMD_{demo}}(\mathcal{P})-\mathsf{MMD_{swap}}(\mathcal{P})$). For the MMD discrepancy measurement, we use chains that were optimized with short bursts to include more majority-Black districts. We say that a chain shows signs of convergence after a certain number of steps when $\hat{R}<1.01$ and $\texttt{ESS}>400$.
All geographies except for the Wyoming state lower house ($\hat{R}=1.06$, $\texttt{ESS}=43.3$) showed signs of convergence after 1,000,000 steps when measuring population balance. Seven geographies did not show signs of converging after 1,000,000 steps under the MMD discrepancy measurement. All six of these geographies have relatively few majority-Black districts; the highest number of majority-Black districts found in any chain from these geographies was three.
@@
OH 1.00008 38567.1 1.00784 1555.93 1.00003 158958 1.01815 205.732
OK 1.00026 24267.1 1.00641 434.785 1.00004 73844.2 1.00066 9354.15
OR 1.00012 50909.7 1.00002 142515
PA 1.00043 14389.1 1.00824 847.099 1.00006 101015 1.00063 4371.36
RI 1.00013 15854.8 1.00009 58597.5
SC 1.0002 19387.5 1.00863 793.865 1.00007 87268.3 1.00558 1167.13
SD 1.00012 46681.3 1.00006 55289.7
TN 1.00017 33200.2 1.0081 491.748 1.00001 156826 1.00449 648.003
TX 1.00012 25229.2 1.00654 1943.14 1 93990.8
UT 1.00002 27696.1 1 131608
VT 1.00001 272274
VA 1.00006 34841.9 1.00915 1230.04 1.00001 115771 1.0045 770.027
WA 1.00002 78639.5 1.00012 79849.2
WV 1.00003 34632 1.00001 244574
WI 1.00006 31502.7 1.00152 3732.32 1.00001 143569 1.01361 323.169
WY 1.05908 43.2695 1.00033 28393
Table 3: Convergence tests for state legislative district ensembles. Cells that are highlighted in red represent chains that do not meet the threshold of $\hat{R}\leq 1.01$, $\mathsf{ESS}\geq 400$, suggesting possible non-convergence. Cells that are highlighted in gray represent ensembles for which the $\hat{R}$ computation is undefined, either because no plans were found that were balanced in population in our experimental setup, or because every plan within an ensemble contained the same number of MMDs. Cells that are highlighted in black represent legislative geographies that do not exist.
Robustness of Critical Offset Computation
In Table 4 we compute the critical offset for state lower house geographies by repeatedly running chains with a higher offset $\Delta$ from the base acceptable $\tau$. In our case, we default to $\tau=5\%$, and $\Delta$ is increased in increments of 0.5%. We repeat this computation ten times for each geography to observe the variation in critical offset between runs. In 17 of the 45 geographies, every test yielded the same critical offset. The largest deviation we observed between the smallest and largest critical offset for a geography was 0.3% for the Wyoming state lower house, with all other geographies having deviations no greater than 0.2%.
State Mean $\Delta$ StDev
AL 0.200 0.000
AK 0.420 0.024
AZ 0.000 0.000
AR 0.250 0.000
CA 0.050 0.000
CO 0.115 0.023
CT 0.495 0.027
DE 0.250 0.000
FL 0.100 0.000
GA 0.245 0.015
@@
OH 0.075 0.025
OK 0.300 0.000
OR 0.145 0.015
PA 0.160 0.020
RI 0.825 0.025
SC 0.235 0.023
SD 0.155 0.015
TN 0.110 0.020
TX 0.100 0.000
UT 0.205 0.015
VT
VA 0.145 0.015
WA 0.060 0.020
WV 0.165 0.023
WI 0.150 0.000
WY 0.585 0.081
Table 4: Mean and standard deviation of critical offset ($\Delta$) computed for each state lower house plan. For each plan, starting at $\Delta=0$, we sample an ensemble using a population tolerance threshold in demo of $5\%-\Delta$, and compute the percentage of plans that exceed a 5% population tolerance. If this percentage is under 2%, the critical offset is $\Delta$, otherwise we increase $\Delta$ by 0.05% and repeat the procedure. The average and standard deviation are computed over 10 experimental runs. Cells that are highlighted in gray represent states where no plans were found that were balanced in population in our experimental setup. Cells that are highlighted in black represent legislative geographies that do not exist.
@@
All results in the paper were reported for chains using the census block group as the base unit of district composition. In Table 5 we report results for our critical offset computations using chains that are composed districts from census blocks, the smallest unit of geography defined in the census hierarchy which is commonly used as the smallest geographic unit in redistricting. Due to computational constraints, we only report results for three legislative geographies: RI state lower house, RI state upper house, and CT state lower house. In all three geographies, we found that using blocks instead of block groups led to an increase in both no-offset discrepancy rate and critical offset.
Block Group Block
@@
RI SH 0.767 1.0 0.914 ¿1.0
RI SS 0.217 0.3 0.405 0.55
CT SH 0.826 0.45 0.976 0.7
Table 5: Comparison between no offset discrepancy rate and critical offset for three geographies using either block groups or blocks as the smallest unit of district construction.
Robustness to District Type We provide metrics from chains run on Congressional districting plans for each state that was apportioned more than one seat in the House of Representatives after the 2010 census. While there is no population deviation that is generally considered de minimis in Congressional redistricting, we adopt the same $\tau=5\%$ used in our other experiments for the sake of consistency. We find that in every state, the critical offset is 0%, meaning that fewer than 2% of plans in each ensemble measured a population deviation in excess of $\tau$ under $\mathop{\sf SWAP}$. Similarly, in every ensemble we found that there existed plans which maximized the number of majority-Black districts in $\mathop{\sf DEMO}$ while having an MMD discrepancy of 0.
@@
State $\Delta$ DR ($\Delta=0$) Mean Discrep. Discrep. Rate Max Agreement
AL 0.0 0.00018 0.01519 0.01519
@@
Table 6: Metrics for ensembles of Congressional district plans. States highlighted in black were apportioned a single congressional representative after the 2010 census, and states highlighted in gray had formatting issues in the underlying boundary data preventing ensembles from being generated. For each ensemble sampled with standard constraints, we calculate the critical offset ($\Delta$) and the no-offset discrepancy rate (i.e. the proportion of plans exceeding a 5% population balance threshold under $\mathsf{SWAP}$). For each ensemble sampled using short burst optimization, we calculate the mean MMD discrepancy, the discrepancy rate (i.e. the proportion of plans with a non-zero MMD discrepancy), and whether there exist plans in the ensemble that contain a maximal number of MMDs under $\mathsf{DEMO}$ and have an MMD discrepancy of zero.
@@
We compare our results from Section 5 to state lower house chains optimized to increase the number of majority-Hispanic districts in a given plan, rather than majority-Black districts. Among the ensembles where majority-Hispanic districts were found, all had no-offset discrepancy rates of less than 50%, whereas the Mississippi state house had a no-offset discrepancy rate of 63.45% when optimized to promote majority-Black districts. Only in the Texas state lower house were there no plans found that were maximal in $\mathop{\sf DEMO}$ while having 0 MMD discrepancy.
Short Bursts
State Mean Discrep. Discrep. Rate Max Agreement
@@
Table 7: Metrics for ensembles optimized to increase the number of majority-Hispanic districts. We sampled an ensemble of 100,000 plans for each state lower house geography using short burst optimization, calculating the mean MMD discrepancy, the discrepancy rate (i.e. the proportion of plans with a non-zero MMD discrepancy), and whether there exist plans in the ensemble that contain a maximal number of MMDs under $\mathsf{DEMO}$ and have an MMD discrepancy of zero.
Appendix B Distribution of BVAP Population in Standard and Optimized Ensembles.

### v2

@@ Acknowledgements
We thank Peter Rock and Conlan Olson for providing data for our experiments, as well as Micah Altman, John Abowd, and Moon Duchin for valuable feedback. The authors were supported by the DARPA SIEVE program under Agreement No. HR00112020021. Preliminary work on this project was supported by the Boston University Center for Antiracist Research.
Appendix A Additional Results
@@
To provide evidence of the convergence of our chains, we report the Gelman-Rubin split-$\hat{R}$ metric and the effective sample size (ESS) for each legislative geography we study in Table 3. Traditionally, an $\hat{R}$ of below 1.1 has been accepted as evidence of convergence, although more recent work recommends that using a threshold of 1.01 for $\hat{R}$ along with a rank-normalized ESS of at least 400 for more robust results. We test for the convergence of our measurement of population balance (i.e. whether our estimate of the mean of the function $f(\mathcal{P})=\mathds{1}\left[\mathsf{dev_{swap}}(\mathcal{P})>0.1\right]$ has converged, where $\mathcal{P}$ is a plan sampled from the stationary distribution of our Markov chain) and our measurement of MMD discrepancy (i.e. $f(\mathcal{P})=\mathsf{MMD_{demo}}(\mathcal{P})-\mathsf{MMD_{swap}}(\mathcal{P})$). For our population balance measurment, we tested convergence using 4 chains of length 1,000,000, subsampling every $10^{th}$ plan to yield an ensemble of 100,000 plans. For the MMD discrepancy measurement, we tested convergence using the 100 ensembles that we generated for each legislative geography as described in Section 6. We say that a chain shows signs of convergence after a certain number of steps when $\hat{R}<1.01$ and $\texttt{ESS}>400$.
All geographies except for the Alabama, Connecticut, West Virginia, and Wyoming state lower houses showed signs of convergence after 1,000,000 steps when measuring population balance. Many geographies did not show signs of converging after 100,000 steps under the MMD discrepancy measurement. However, given our method for creating ensembles with lots of MMDs (sampling maximal plans from each of the 100 chains that we ran), the convergence of the individual chains doesn’t have a straightforward relationship to that of the final ensemble. Further work is necessary to characterize the distribution of plans that maximize majority-minority districts.
@@
OH 1 40349.8 1.04475 1415.6 0.999999 143535 1.01950 3294.4
OK 1.00027 22396.2 1.12866 488.9 1.00013 66426.5 1.00780 12631.9
OR 1.00015 50394.5 1.00000 137612.0
PA 1.00016 16996.6 1.06915 892.7 1.00007 93319.3 1.01360 6143.2
RI 1.00014 14566.3 1.00006 58658.8
SC 1.00032 16261.9 1.08331 769.2 1.00001 80159.9 1.04219 1449.7
SD 1.00012 43806.1 1.00004 50200.9
TN 1.00011 19834.9 1.11814 543.6 1.00001 140019 1.05512 1069.4
TX 1.00011 23232.5 1.01942 4100.9 1.00003 190915
UT 1.0002 29949.4 1.00001 110132
VT 1.00001 171445
VA 1.00014 23230.8 1.04422 1397.1 1 107197 1.07987 755.4
WA 1.00004 77180.7 1.00003 73462.3
WV 1.20887 13.0 1 237573
WI 1.00003 29893.9 1.02260 3215.7 1.00004 134723 1.10688 580.7
WY 1.08295 31.8 1.00018 37894
Table 3: Convergence tests for state legislative district ensembles. Cells that are highlighted in red used block groups as the base unit of geography due, all others used precincts. Cells that are highlighted in gray represent ensembles for which the $\hat{R}$ computation is undefined, either because no plans were found that were balanced in population in our experimental setup, or because every plan within an ensemble contained the same number of MMDs. Cells that are highlighted in black represent legislative geographies that do not exist.
Population Balance and Critical Offsets
In Table 4 we compute no offset discrepancy rates and the critical offsets for state lower and upper house geographies. The critical offsets are computed by repeatedly running chains with a higher offset $\Delta$ from the base acceptable $\tau$. For each geography, starting at $\Delta=0$, we sample an ensemble using a population tolerance threshold in demo of $5\%-\Delta$, and check whether any plan exceeds a 5% population tolerance. If this there does not exist such a plan in the ensemble, the critical offset is $\Delta$, otherwise we increase $\Delta$ by 0.05% and repeat the procedure. The average and standard deviation are computed over 10 experimental runs. We default to $\tau=5\%$, and $\Delta$ is increased in increments of 0.05%. We repeat this computation ten times for each geography to observe the variation in critical offset between runs. The largest deviation we observed between the smallest and largest critical offset for a geography was 0.55% for the Montana state lower house, with all other geographies having deviations no greater than 0.5%. We also report what are synthetic model would predict for no offset discrepancy rate and critical offset for each geography, as described in Section 5.2.
State Lower House State Upper House
Discrepancy ($\tau=5\%$) Critical Offset ($\Delta$) Discrepancy ($\tau=5\%$) Critical Offset ($\Delta$)
State Observed Predicted Mean $\Delta$ (StDev) Predicted $\Delta$ Observed Predicted Mean $\Delta$ (StDev) Predicted $\Delta$
AL 63.280% 68.802% 0.535% (0.045%) 0.540% 19.088% 17.617% 0.270% (0.024%) 0.310%
AK 65.634% 70.735% 1.580% (0.100%) 1.670% 26.436% 29.790% 0.940% (0.044%) 0.890%
AZ 13.917% 13.464% 0.235% (0.039%) 0.265% 14.068% 13.366% 0.240% (0.020%) 0.255%
AR 87.448% 86.045% 0.955% (0.057%) 1.090% 29.745% 26.677% 0.435% (0.032%) 0.460%
CA 23.843% 20.163% 0.155% (0.015%) 0.150% 8.144% 6.425% 0.100% (0.000%) 0.100%
CO 53.002% 51.869% 0.550% (0.059%) 0.600% 24.416% 22.247% 0.350% (0.032%) 0.385%
CT 89.200% 96.267% 0.717% (0.062%) 1.215% 22.958% 24.228% 0.370% (0.040%) 0.400%
DE 58.165% 63.885% 0.970% (0.060%) 1.320% 24.496% 26.760% 0.670% (0.033%) 0.755%
FL 58.247% 53.510% 0.350% (0.045%) 0.350% 12.441% 10.501% 0.140% (0.020%) 0.150%
GA 90.185% 90.446% 0.600% (0.050%) 0.735% 26.967% 26.758% 0.275% (0.025%) 0.310%
@@
OH 52.874% 47.655% 0.325% (0.040%) 0.365% 12.567% 9.318% 0.155% (0.015%) 0.155%
OK 84.469% 84.903% 0.970% (0.071%) 1.050% 40.591% 40.346% 0.540% (0.037%) 0.570%
OR 41.871% 43.524% 0.530% (0.051%) 0.535% 15.693% 15.504% 0.300% (0.022%) 0.305%
PA 85.595% 86.451% 0.545% (0.057%) 0.550% 18.750% 15.653% 0.170% (0.024%) 0.195%
RI 86.942% 90.707% 1.795% (0.117%) 1.730% 48.021% 52.211% 1.225% (0.093%) 1.040%
SC 84.431% 85.243% 0.690% (0.049%) 0.845% 29.123% 28.846% 0.350% (0.032%) 0.400%
SD 54.947% 54.113% 1.020% (0.064%) 1.110% 46.376% 50.573% 0.995% (0.072%) 1.045%
TN 64.002% 62.194% 0.485% (0.071%) 0.540% 14.440% 13.176% 0.225% (0.025%) 0.230%
TX 62.132% 58.322% 0.310% (0.030%) 0.350% 8.014% 5.564% 0.105% (0.015%) 0.100%
UT 75.406% 74.807% 0.880% (0.051%) 1.005% 25.152% 23.305% 0.425% (0.040%) 0.460%
VT 4.730% 5.160% 0.290% (0.020%) 0.200%
VA 57.649% 57.934% 0.405% (0.042%) 0.485% 16.420% 15.990% 0.225% (0.025%) 0.240%
WA 37.025% 32.980% 0.400% (0.039%) 0.440% 38.125% 33.107% 0.400% (0.000%) 0.445%
WV 66.127% 66.311% 0.750% (0.047%) 0.880% 8.815% 8.598% 0.250% (0.032%) 0.265%
WI 65.347% 64.443% 0.555% (0.052%) 0.570% 16.543% 13.467% 0.230% (0.033%) 0.225%
WY 100.000% 94.393% 2.175% (0.144%) 2.645% 56.729% 57.540% 1.440% (0.073%) 1.485%
Table 4: Empirical and Predicted Discrepancies and Critical Offsets for 93 Legislative Geographies. Critical offsets are computed as described in Appendix A (averaged over 10 runs). Columns labeled “Predicted” are computed by sampling discrepancies (with and without offsets) from the probabilistic model described in Section 5.2 using the empirical means and standard deviations of the corresponding legislative redistricting ensembles. Cells that are highlighted in gray represent states where no plans were found that were balanced in population in our experimental setup. Cells that are highlighted in red had precinct data that was incompatible with our experimental setup (either due to large precincts or a lack of complete precinct data) and were computed using block groups as the building blocks of districts instead. Cells that are highlighted in black represent legislative geographies that do not exist.
@@
All results in the paper were reported for chains using either precincts or census block groups as the base unit of district composition. In Table 5 we report results for our critical offset computations using chains that are composed from census blocks, the smallest unit of geography defined in the census hierarchy which is sometimes used as the base unit in redistricting. Due to computational constraints, we only report results for three legislative geographies: RI state lower house, RI state upper house, and CT state lower house. In all three geographies, we found that using blocks instead of block groups led to an increase in both no-offset discrepancy rate and critical offset. See Section 7 for additional discussion about the possible effects of the choice of geounit.
Block Group/Precinct Block
@@
CT SH 89.20% 0.72% 98.58% >>2.0%
RI SH 86.94% 1.80% 97.62% >>2.0%
RI SS 48.02% 1.23% 66.14% 1.60%
Table 5: Comparison between no offset discrepancy rate and critical offset for three geographies using precincts (Connecticut) or block groups (for Rhode Island legislative geographies) versus blocks as the smallest unit of district construction. We halted the computation of critical offsets at 2.0%.
Robustness to District Type
We provide metrics from chains run on Congressional districting plans for each state that was apportioned more than one seat in the House of Representatives after the 2010 census. While there is no population deviation that is generally considered de minimis in Congressional redistricting, we adopt the same $\tau=5\%$ used in our other experiments for the sake of consistency. We find that in every state, the average critical offset over 10 runs, as defined in Section 5, is less than 0.2%.
@@
State $\Delta$ DR ($\Delta=0$) Mean Discrep. Discrep. Rate
AL 0.10% 1.67% 0.18 17.20%
@@
Table 6: Metrics for ensembles of Congressional district plans. Cells highlighted in black represent states that were apportioned a single congressional representative after the 2010 census, cells highlighted in gray represent geographies for which no majority-Black districts were found by our chains, and cells highlighted in red represent chains that were run using block groups as the base unit of geography due to missing precinct data. For each ensemble sampled with standard constraints, we calculate the critical offset ($\Delta$) and the no-offset discrepancy rate (i.e. the proportion of plans exceeding a 5% population balance threshold under $\mathsf{SWAP}$). Critical offsets are averaged over 10 runs. For each ensemble sampled using short burst optimization, we calculate the mean MMD discrepancy and the discrepancy rate (i.e. the proportion of plans with a non-zero MMD discrepancy).
@@
We compare our results from Section 5 to state lower house chains optimized to increase the number of majority-Hispanic districts in a given plan, rather than majority-Black districts. Each optimized ensemble is constructed by running 100 ReCom chains with short bursts optimization to generate plans with many majority Hispanic districts, and sampling 10 plans from each chain with the highest number of majority Hispanic districts to be included in the ensemble. Each chain uses a population tolerance threshold of $5\%-\Delta$, where $\Delta$ is the critical offset for that legislative geography, as computed in Section 5. The magnitudes of discrepancies we observed for majority-Hispanic districts are similar in magnitude to those of majority-Black districts. For example, the largest mean discrepancy we observed for majority-Hispanic districts was 1.093 in the New Mexico State Lower House, while the larges we observed for majority-Black districts was 1.431 in the Mississippi State Lower House.
State Lower House State Upper House
State Non-Zero Discrep. Rate Mean Discrep. Non-Zero Discrep. Rate Mean Discrep.
@@
Table 7: Metrics for ensembles optimized to increase the number of majority-Hispanic districts. We show the mean majority-Hispanic district discrepancy and the discrepancy rate (i.e. the proportion of plans with a non-zero MMD discrepancy). Gray cells represent geographies for which our ensembles did not contain any majority-Hispanic districts in either DEMO or SWAP. Black cells represent legislative geographies that do not exist.
MMD Discrepancies Across Geographies
Table 8 contains the mean MMD discrepancy and non-zero discrepancy rate for optimized ensembles across 52 legislative geographies. Each optimized ensemble is constructed by running 100 ReCom chains with short bursts optimization to generate plans with many majority-Black districts, and sampling 10 plans from each chain with the highest number of majority-Black districts to be included in the ensemble. Each chain uses a population tolerance threshold of $5\%-\Delta$, where $\Delta$ is the critical offset for that legislative geography, as computed in Section 5. The legislative geographies included are those for which our chains found at least one plan with at least one majority-minority district.
Short Bursts
State Non-Zero Discrep. Rate Mean Discrep. Non-Zero Discrep. Rate Mean Discrep.
AL 57.5% 0.822 25.3% 0.293
AK
AZ
AR 37.2% 0.470 54.2% 0.715
CA 0.9% 0.009
CO
CT 11.6% 0.126 6.1% 0.061
DE 11.7% 0.120 4.3% 0.045
FL 23.4% 0.256 7.6% 0.076
GA 65.6% 1.027 34.7% 0.433
HI
ID
IL 23.3% 0.270 8.9% 0.095
IN 31.5% 0.364 28.2% 0.291
IA
KS 39.2% 0.415
KY 2.4% 0.024
LA 68.3% 1.122 43.7% 0.553
ME
MD 17.6% 0.191 16.6% 0.181
MA 17.6% 0.181 5.4% 0.054
MI 11.3% 0.120 6.4% 0.067
MN
MS 78.9% 1.431 53.7% 0.747
MO 32.9% 0.405 58.3% 0.790
MT
NE 38.9% 0.389
NV
NH
NJ 11.0% 0.114 10.1% 0.107
NM
NY 42.7% 0.533 17.5% 0.187
NC 49.2% 0.649 27.8% 0.315
ND
OH 24.8% 0.289 24.3% 0.261
OK 34.7% 0.404 15.0% 0.150
OR
PA 20.2% 0.220 14.3% 0.149
RI
SC 71.3% 1.203 59.3% 0.847
SD
TN 30.5% 0.325 26.0% 0.291
TX 9.2% 0.096
UT
VT
VA 31.6% 0.367 53.1% 0.724
WA
WV
WI 20.3% 0.218 46.3% 0.493
WY
Table 8: Mean majority-Black district discrepancy and non-zero discrepancy rate for short-burst optimized ensembles across 52 state legislative geographies. Gray cells represent geographies for which our ensembles did not contain any majority-Black districts in either DEMO or SWAP. Black cells represent legislative geographies that do not exist. Red cells represent chains that were run using block groups as the base unit of geography due to missing precinct data.
Distribution of BVAP Population in Standard and Optimized Ensembles.
@@
Appendix B Towards Disentangling the Impact of the 2010 and 2020 Disclosure Avoidance Systems
We would like to understand the impact of the 2020 DAS on MMD discrepancies. Unfortunately, comparing DEMO and SWAP reflects the combined effects of the 2010 DAS and the 2020 DAS. This is because DEMO is the result of applying the 2020 DAS to the confidential 2010 CEF data, while SWAP is the result of applying the 2010 DAS to the confidential 2010 CEF data.
One way to try to isolate the effect of the 2020 DAS would be to run it ourselves using reconstructed 2010 Census microdata derived from SWAP (as in [12]). But running the 2020 DAS reliably is difficult [4] and beyond the scope of the present work.
Instead, we try to isolate the effect of the 2010 DAS using a dataset provided to us by the authors of [4] (personal communication, September 2025). The dataset is the result of applying an implementation of swapping (designed to approximate the 2010 DAS as closely as possible using only publicly-available information) to reconstructed 2010 Census microdata for Georgia, as described in [4]. Hence, this dataset—which we call SWAP-SWAP—is the result of applying the swapping-based 2010 DAS to the 2010 CEF, then reconstructing it and applying an independent implementation of swapping.
We generate a new ensembles of plans using the SWAP-SWAP and DEMO datasets (using short-bursts optimization for $\mathsf{ReCom}$ with 100 chains in parallel). For each ensemble, we measure the MMD discrepancies relative to SWAP. (For this analysis only, we count majority-Black districts using the total population rather than the voting age population, because voting age data was not included in the SWAP-SWAP dataset.) The results are plotted in Figure 10. The similarities are clear. In both ensembles, the number of MMDs according the dataset used to generate the plans (DEMO, SWAP-SWAP) tends to be greater than the number of MMDs according to the unseen dataset (SWAP). Moreover, the bias grows with the number of MMDs.
Going a step further, SWAP-SWAP population data makes it possible to estimate what Black population margins might look like in the CEF. This holds under a strong assumption on the fidelity of the SWAP-SWAP data. [^14] It appears difficult or impossible to validate this assumption for the same reasons that measuring the effects the 2010 swapping-based DAS is difficult: the swapping algorithm and the CEF input data is secret, known only to the Census Bureau. Namely, the analysis below requires the differences between tabulations in SWAP-SWAP and SWAP be distributed similarly to the differences between tabulations in SWAP and the 2010 CEF. Adapting the notation introduced in Section 6.3,
$B_{\mathsf{swap\text{-}swap}}(D)-B_{\mathsf{swap}}(D)\approx B_{\mathsf{swap}}(D)-B_{\mathsf{cef}}(D),$ (3)
where $B_{\mathsf{data}}(D)$ is the Black population margin (i.e., $\mathsf{pop}^{\mathsf{Black}}_{\mathsf{data}}(D)-0.5\cdot\mathsf{pop}^{\mathsf{Black}}_{\mathsf{dara}}(D)$) in district $D$ as measured in dataset DATA. This differs slightly from the definition of $B_{\mathsf{data}}(D)$ in Section 6.3, which considers voting age population rather than total population. This is due to the lack of voting age information in SWAP-SWAP, as mentioned above.
Assuming the above, we can approximate Black population margins in the 2010 CEF directly from the margins we measured using SWAP and SWAP-SWAP:
$B_{\mathsf{cef}}(D)$ $=B_{\mathsf{swap}}(D)+\bigl(B_{\mathsf{cef}}(D)-B_{\mathsf{swap}}(D)\bigr)$ (4)
$\approx B_{\mathsf{swap}}(D)+\bigl(B_{\mathsf{swap}}(D)-B_{\mathsf{swap\text{-}swap}}(D)\bigr)$ (5)
$=2B_{\mathsf{swap}}(D)-B_{\mathsf{swap\text{-}swap}}(D)$ (6)
Using these estimates, we plot MMD discrepancies between DEMO and the above estimate of CEF in Figure 11. Discrepancies between DEMO and CEF are qualitatively similar to discrepancies between DEMO and SWAP, although slightly smaller in magnitude. This result suggests that our observations in Section 6 would have been similar had we been using the 2010 CEF rather than SWAP.
To reiterate, the analysis above assumes that the implementation of swapping used to create our SWAP-SWAP dataset closely approximates the 2010 DAS. More work would be needed to fully disentangle the effects of the 2020 and 2010 disclosure avoidance systems, even setting aside the extent to which comparing SWAP-SWAP to SWAP is a faithful representation of the effects of the 2010 DAS. For example, it is not clear whether the similarities between the to plots in Figure 10 mean that the effect of the 2020 DAS is smaller, larger, or similar to the effects of swapping. Instead, we remark only that any differences in the MMD discrepancies arising from the two systems appear more likely differences in degree than in kind.
Figure 10: MMD discrepancies in SWAP-SWAP. Distribution of MMD discrepancies (relative to SWAP) for Georgia state house using the DEMO and SWAP-SWAP datasets. The number of MMDs on the horizontal axis is measured using the dataset used to generate the ensembles (DEMO or SWAP-SWAP), and the percent of plans normalized to 100% for each bin is displayed on the vertical axis. We generate ensembles using the method described in Section 6.1, running 100 chains for each dataset.
Figure 11: MMD discrepancies in Reconstructed CEF Data. Distribution of MMD discrepancies for an ensemble using short bursts optimization to sample plans with many MMDs in DEMO. On the left, we display the MMD discrepancy between DEMO and SWAP. On the right, we display the MMD discrepancy between DEMO and our estimation of the 2010 CEF, as described in Appendix B.